OpenStudy (anonymous):

1. What is the equation of a line in slope-intercept form that has slope -2 and passes through the point (1, 2)?

OpenStudy (mathmale):

You are given the slope of this line, along with one point thru which the line passes. To me this is an invitation to use the point-slope form of the equation of a straight line. Would you do that, please? Once you have that equation, you can easily re-write it in slope-intercept form. Alternatively, Frankie, jump directly into the slope-intercept form: y = m*x + b We know one point on the line: (1,2), along with the slope, m=-2. Substitute this info into y = m*x + b and then solve for b. Finally, write your equation / formula in the form y = m*x + b, using the value of b that you've just calculated.

OpenStudy (anonymous):

I got this on my own but thanks

OpenStudy (mathmale):

good for you!